For a three sector economy the following is given : C= 30+0.75Y, I = 30 , G=40 . Where C= consumption I = investment and G = government expenditure. Find out the equilibrium output level. Find out the value of investment multiplier.

1
Expert's answer
2021-12-20T10:20:23-0500

i)Equilibrium output level

"Y=C+I+G"


"Y=30+0.75Y+30+40"


"0.25Y=100"


"Y=\frac{100}{0.25}=400"



ii)Investment Multiplier, "k = \dfrac {∆Y}{∆I}"


It is the ratio of the change in national income induced by a $1 change in investment.

Given the consumption function,

"C = 30 + 0.75Y"

0.75 is the marginal propensity to consume (MPC)

Since, MPC + MPS = 1

=> MPS = 1 - MPC

= 1 - 0.75

= 0.25


Now, "K = \dfrac {∆Y}{∆I} = \dfrac {1}{MPS}"


"= \dfrac {1}{1-0.75}"


"=\dfrac {1}{0.25}"


"= \bold {4}"